In each of the following examples, first calculate the PED and then decide if the PED is elastic,

inelastic, perfectly elastic, perfectly inelastic or unitary elastic.

a. The price of a product falls from $8 to $6, causing demand to extend from 10,000 to

12,500.

b. Demand contracts from 500 to 400 when price rises from $40 to $42.

c. Demand extends from 2,000 to 2,800 when price falls from $20 to $18.

d. Price rises from $15 to $30 but demand stays unchanged at 5,000.

e. An increase in price from $80 to $90 and as a consequence, a decrease in quantity

demanded from 400 to 300. 



1
Expert's answer
2022-03-16T09:44:05-0400

a) PED= "\frac {\Delta Q}{\Delta P}× \frac{P}{Q}"

"\frac {(12500-10000)}{(8-6)}× \frac{8}{10000}= -1"

This is unitary elastic

b) "PED=\frac {\Delta Q}{\Delta P}× \frac{P}{Q}"

"\frac {-100}{2}× \frac{40}{500}= -4"

This is perfectly elastic

c) "\frac {800}{-2}× \frac{20}{2000}= -4"

This is perfectly elastic

d) "\frac {0}{15}× \frac{15}{5000}= 0"

This is perfectly inelastic

e) "\frac {-100}{10}× \frac{80}{400}= -8"

This is perfectly elastic